# Don’t Forget Its Variance! The Minimum Path Variance Principle for Accurate and Stable Score-Based Density Ratio Estimation

### v1

@@ Abstract
Score-based methods have emerged as a powerful framework for density ratio estimation (DRE), but they face an important paradox in that, while theoretically path-independent, their practical performance depends critically on the chosen path schedule. We resolve this issue by proving that tractable training objectives differ from the ideal, ground-truth objective by a crucial, overlooked term: the path variance of the time score. To address this, we propose MinPV (Minimum Path Variance) Principle, which introduces a principled heuristic to minimize the overlooked path variance. Our key contribution is the derivation of a closed-form expression for the variance, turning an intractable problem into a tractable optimization. By parameterizing the path with a flexible Kumaraswamy Mixture Model, our method learns a data-adaptive, low-variance path without heuristic selection. This principled optimization of the complete objective yields more accurate and stable estimators, establishing new state-of-the-art results on challenging benchmarks.
@@ 1 Introduction
The key theoretical result of these methods is elegant: The exact log-density ratio is recovered by integrating the time score along any smooth path (Choi et al., 2022), making the result path-invariant in theory. However, a striking paradox emerges in practice: When this time score is approximated by a neural network, the final performance of the estimator exhibits a significant dependence on the chosen path, as shown in our preliminary experiments (see Fig. 1(a)).
We trace the problem to a subtle but crucial gap between the ideal objective, which minimizes the true mean-squared error to the ground-truth time score, and the tractable score-matching loss used in practice. The two objectives differ by a path-dependent term, ignored when the path is fixed since it becomes a constant (Choi et al., 2022). However, our analysis reveals that this term is in fact the dominant factor driving performance differences across paths, akin to the impact of noise schedules in generative modeling (Kingma et al., 2021; Guo and Schwing, 2025; Strasman et al., 2025).
@@
Our main contribution is to close this gap. We prove that the missing term is precisely the path variance of the ground-truth time score, a quantity that must be minimized and cannot be treated as a constant (illustrated in Figs. 1(b) and 1(c) and proved in Sec. 4.1). Building on this insight, we propose MinPV (Minimum Path Variance) Principle, the first framework that directly optimize the complete, ideal objective. Our contributions are threefold:

### v2

# Don’t Forget Its Variance! The Minimum Path Variance Principle for Accurate and Stable Score-Based Models

@@ Abstract
Score-based methods are powerful across machine learning, but they face a paradox: theoretically path-independent, yet practically path-dependent. We resolve this by proving that practical training objectives differ from the ideal, ground-truth objective by a crucial, overlooked term: the path variance of the score function. We propose the MinPV (Minimum Path Variance) Principle to minimize this path variance. Our key contribution is deriving a closed-form expression for the variance, making optimization tractable. By parameterizing the path with a flexible Kumaraswamy Mixture Model, our method learns data-adaptive, low-variance paths without heuristic manual selection. This principled optimization of the complete objective yields more accurate and stable estimators, establishing new state-of-the-art results on challenging benchmarks and providing a general framework for optimizing score-based interpolation.
@@ 1 Introduction
Theoretically, these methods are path-invariant: integrating the score along any smooth path recovers the exact target (Choi et al., 2022). However, a practical paradox emerges in practice: with neural network approximation, performance becomes significantly path-dependent, as shown in our preliminary experiments (see Fig. 1(a)). This sensitivity mirrors challenges in broader probabilistic inference and generative modeling (Sahoo et al., 2024; Strasman et al., 2025).
We trace this paradox to a subtle but crucial gap between the ideal objective, which minimizes the true mean-squared error to the ground-truth score, and the tractable score-matching loss used in practice. The two objectives differ by a path-dependent term that is typically ignored as a constant when the path is fixed (Choi et al., 2022). Our analysis reveals that this term is, in fact, the dominant factor driving performance differences across paths, akin to the impact of noise schedules in generative modeling (Kingma et al., 2021; Guo and Schwing, 2025; Strasman et al., 2025).
@@
Our main contribution is to close this gap. We prove that the missing term is precisely the path variance of the ground-truth score, a quantity that must be minimized and cannot be treated as a constant (illustrated in Figs. 1(b) and 1(c) and proved in Sec. 4.1). Building on this insight, we propose MinPV (Minimum Path Variance) Principle, the first framework that directly optimize the complete, ideal objective. Our contributions are threefold: